Problem 4 (Conditional Expectation and Variance). Suppose the joint distri- bution of (X, Y) is given by the following contingency (row represents x) table 20 points (x,y) 2 4 6 1 0.3 0 0.1 2 0 0.2 0 3 0.1 0 0.3 A) Compute the marginal distributions of X and Y B) Are X and Y independent? Explain. C) Find the conditional distribution of Y given X -1 D) Compute E[Y|X 1] E) Compute EY|X= 2] F Compute E[exp(X)Y|x 2
Answer #1
x,y 2 4 6 total
1 0.3 0 0.1 0.4
2 0 0.2 0 0.2
3 0.1 0 0.3 0.4
total 0.4 0.2 0.4 1
A) marginal distribution of Y
Y 2 4 6 total
P(Y=y) 0.4 0.2 0.4 1
marginal distribution of X
X 1 2 3 total
P(X=x) 0.4 0.2 0.4 1

B)

X and Y are independent if for each possible values of X and Y following is true:

Here P(Y=2) = 0.4

P(X=1) = 0.4

and P(X=1, Y=2) = 0.3

P(X=1)P(Y=2) = 0.4*0.4 =0.16

Since P(X=1)P(Y=2) is not equal to P(X=1, Y=2) so X and Y are not independent.

c)

P(Y|X=1)=P(X=1,Y)/P(X=1)

Y 2 4 6 total
P(Y|X=1) 0.3/0.4=0.75 0 0.1/0.4=0.25 1

d)

Y 2 4 6 total
P(Y|X=1) 0.75 0 0.25 1
Y*P(Y|X=1)= 1.5 0 1.5 3

E[Y|X=1] = ΣY*P(Y|X=1) = 3

e)

P(X=2)=0.2

Y 2 4 6 total
P(Y|X=2) 0/0.2=0 0.2/0.2=1 0/0.2=0 1
Y*P(Y|X=2)= 0 4 0 4

E[Y|X=2] = ΣY*P(Y|X=2) = 4

-----------------------------------------------------------------------------------------------------------------

only four parts can be answered per post as HOMEWORKLIB RULES answering guidelines.